Write code that when you are given the range and probability distribution of a random variable X. (i.e X ="the number of heads showing on 2 flipped fair coins":

range = (0,1,2)

probability distribution: [.25, .5, .25]

Such that the code returns the expected value, standard deviation, and variance of the random variable . thanks!
Answer #1

Solution :

expected value = \sum X * P(X)

= 0 *0.25 + 1 * 0.5 + 2 *0.25  

= ( 0+ 0.5 + 0.5 )

= 1

Standard deviation = \sigma

==\sqrt\sumX 2 * P(X) - \mu 2

=  \sqrt[ 02   * 0.25+ 12   *0.5 + 22 *0.25 ] - (1)2

= \sqrt [( + 0+ 0.5+ 1) - 1]

= \sqrt 1.5 - 1

=\sqrt0.5 =0.7071

variance =\sumX 2 * P(X) - \mu 2

=[ 02   * 0.25+ 12   *0.5 + 22 *0.25 ] - (1)2

=  [( + 0+ 0.5+ 1) - 1]

= 1.5 - 1

=0.5